OpenStudy (anonymous):

Seiji and Gavin both worked hard over the summer. Together, they earned a total of $425. Gavin earned $25 more than Seiji. How much did each of them earn?

OpenStudy (anonymous):

Answer: w = width w + 5 = length P = perimeter = 2w+2w+10 = 4w + 10 4w+10 > 30 4w > 20 w > 5 x = original price. 85x = 45 x = 45/.85 = $52.94 S = Seiji G = Gavin S + G = 425 G = S + 25 So S + S +25 = 425 2S +25 = 425 2S = 400 S = 200 G = 225

OpenStudy (anonymous):

I hope that might have helped